OpenStudy (anonymous):

The length of the base of a triangle is twice its height. If the area of the triangle is 64 sq kilo, find the height. (Theres a triangle and the height is x the width of 2x)

OpenStudy (anonymous):

Height = x base = 2x Area of a traingle = 1/2 base x height So, 64 = (1/2)(x)(2x) 64 = x^2 x = 8 Height = 8 km.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Welcome.